OpenStudy (anonymous):

Convert the linear equation 3x - 2y = 5 to point slope form. Use the point with an x coordinate equal to 1

OpenStudy (lukecrayonz):

Point slope form: y=mx+b where b is your y intercept, and m is your slope. So, solve for y.

OpenStudy (anonymous):

i got y = 3/2 x - 2.5

OpenStudy (anonymous):

what about the last part x coordinate equaling to 0

OpenStudy (anonymous):

to 1 i mean

OpenStudy (lukecrayonz):

Input 1 for x, (you changed your numbers) and then solve for y.

OpenStudy (lukecrayonz):

so (3/2)*1-2.5 equals what

OpenStudy (anonymous):

okay thanks so that will be the final answer

OpenStudy (lukecrayonz):

Yes, but you have to solve the equation in which you should get -1 (3/2)*1=1.5-2.5=-1